Circular No. 7303

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999el IN NGC 6951 C. C. Dudley and J. Fischer, Naval Research Laboratory, Washington,

report near-infrared observations of SN 1999el obtained with the U.S. Naval Observatory 1.55-m telescope (+ ASTROCAM 1024×1024 ALADDIN InSb imager) at Flagstaff on Oct. 26.15 UT: “A K′ position, calibrated from the positions of more than 20 USNO-A2.0 stars within the 6′ field-of-view of the imager, is α = 20^{h}37^{m}17^{s}.72, δ = +66^{o}06′11′′.5 (equinox 2000.0). The K′ and J magnitudes are estimated to be 12.7 and 13.6, respectively (each with an uncertainty of 0.2 mag). Flux calibration is based on the stars in field AS40 reported by Hunt et al. (1998, A.J. 115, 2594), observed within 0.1 airmass of the supernova. The present estimates include contributions from two nearby foreground stars (see IAUC 7288), contributing at the 5–15percent level. Contributions from NGC 6951 appear to be well subtracted by making use of the symmetry of the galaxy in these passbands with respect to rotation by 180^{o} about the nucleus.”

XTE J1859+226 M. Uemura and T. Kato, Kyoto University; and E. Pavlenko, S. Shu-

garov, and M. Mitskevich, Crimean Astrophysical Observatory, communicate: “CCD photometry of the optical counterpart of XTE J1859+226 (IAUC 7274, 7276) with the 0.25-m and 0.38-m telescopes at Kyoto and Crimea revealed likely orbital modulations with a period of 0.2806 day and an amplitude of 0.10 mag. The variations were present on Oct. 20–26, and possibly on Oct. 17, which makes this the earliest detection of such signals among x-ray novae (cf. IAUC 5259, 5692).”

GRB 991105 M. J. S. Smith et al. inform us that the position on IAUC 7301 requires

correction by 4′.4 to α = 12^{h}03^{m}29^{s}, δ = −66^{o}45′24′′ (equinox 2000.0).

SUPERNOVA 1999em IN NGC 1637 CCD magnitudes, unfiltered unless otherwise noted: Nov. 1.000 UT,

13.9 ± 0.2 (K. Sárneczky and L. Kiss, Szeged, Hungary); 1.033, V = 14.32 ± 0.03 (D. Hanžl, Brno, Czech Republic).

1999 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green